OpenStudy (anonymous):

Solve for x as if x was a positive number. x^2 + x = 6 I got 3 or -3 as an answer, but I don't trust myself!

OpenStudy (anonymous):

3^2 + 3 is 12 not 6 2^2 + 2 is 6 however. And since you need a positive answer, you don't need to work out any negative examples.

OpenStudy (anonymous):

x^2+x-6=0

OpenStudy (anonymous):

x^2+3x-2x+6=0 x(x+3)-2(x+3)=0 (x-2)(x+3)=0 x=2,x=-3

OpenStudy (anonymous):

solving x give us 2 and -3